using partial fraction in s domain.
5(5%+4) (5425) 5.(5+6) S
Answer #1

Let- GC8= 5(8²+21) (8²+25) = 5(34+ 29 8²4100) SCS2+62 s3+168 Giso 584 +1455² +500 S3+16S it cara be written as- Ges= 5 st 568مك 2 put s=-1, 505= 17A 4B-C 50 5= 17 (31125 +1B-C B-c =-26-252 - and 2B = -52-5 oroso or 16=-26-25 so -26-25-c= 26.25 so c=0